# Photon rings and shadows of Kerr black holes immersed in a swirling universe

### v1

@@ Abstract
We discuss photon rings around as well as shadows of Kerr black holes immersed in a swirling spacetime. We find that the spin-spin interaction between the angular momentum of the black hole and the swirling of the background leads to new interesting effects as it breaks the symmetry between the upper and lower hemispheres. One of the new features of the spin-spin interaction is the existence of up to three light rings for suitable choices of the angular momentum parameter $a$ and swirling parameter $j$. In comparison to the Schwarzschild black hole immersed in a swirling universe, the light rings typically all possess different radii.
@@ 1 Introduction
Recently, a new exact, rotating solution has been constructed using the Ernst formalism [4]. This solution describes a rotating Kerr black hole in a swirling universe [5]. The swirling universe, i.e. the solution without the black hole, is a vacuum solution equipped with an odd-$\mathbb{Z}_{2}$ symmetry: the upper ($\theta\in[0:\pi/2[$ ) and lower hemispheres ($\theta\in\ ]\pi/2:\pi]$) of the spacetime rotate in opposite directions. The swirling universe can, in fact, be obtained by using the Lie point symmetry through the application of the Ehlers transformation. Note that the geodesic motion can be decoupled and completely integrated in this spacetime [6]. An interesting question is whether this solution is of potential observational relevance, and which new features appear that allow distinction from the “pure” Kerr spacetime. An analysis of accretion discs in this spacetime has already been performed, revealing their bowl-like structure [7]. Recently, the interest in generating techniques has grown again, and different new solutions including the swirling background have been constructed. In particular, external electromagnetic fields were considered in a composition of Harrison and Ehlers’ transformation [8], as well as charged black holes, i.e. Kerr-Newman black holes immersed in the swirling background [9, 10].
@@ 2 The spacetime and its properties
The metric tensor describing the KBHSU can be generated by applying an Ehlers transformation on a Kerr black hole seed solution within the Ernst formalism [4]. In Boyer-Lindquist coordinates, it reads [4] :
@@
$\omega_{0}$ $=$ $2aMr,\ \ \ \quad\omega_{1}=4\cos\theta(Ma^{4}-r^{4}(r-2M)-\Delta a^{2}r-a(r-M)\mathcal{W}),$
$\omega_{2}$ $=$ $2M\left[3ar^{5}-a^{5}(r+2M)+2a^{3}r^{2}(r+3M)-r^{3}(\cos^{2}\theta-6)\mathcal{W}+a^{2}\mathcal{W}\left(\cos^{2}\theta(3r-2M)-6(r-M)\right)\right],$
@@ 2.1 Conical singularity
The spacetime is non-asymptotically flat and has a non-removable conical singularity. We can identify the existence of an axial angular deficit/excess by considering the ratio between the perimeter and radius of a small circle around the rotational axis at both the north and the south poles [4, 9, 10]
@@
for non-vanishing angular momentum parameter $a$ and finite $j$. Thus the SBHSU does not suffer from conical singularities. We note the divergence for the critical value $ajM=\pm 0.25$.
@@ 2.2 Ergoregions
A typical characteristic observed in the swirling family of solutions is the appearance of non-compact ergoregions extending from the equatorial plane to infinity [4, 7, 9, 10]. An ergoregion is defined by $g_{tt}>0$ with boundary at $g_{tt}=0$. For the KBHSU spacetime this reads $\omega^{2}/{\cal F}-\rho^{2}{\cal F}>0$ with boundary at $\omega^{2}=\rho^{2}{\cal F}^{2}$. The angular velocity $\Omega$ of the frame dragging is given by $\Omega=-g_{t\varphi}/g_{\varphi\varphi}=\omega$, which is not constant. We focus our analysis on the region outside the event horizon. The cross section of the ergoregions with the plane $y=0$ ($\varphi=0,\pi$) and their frame dragging directions are shown in Fig. 2 for two values of the angular momentum parameter, $a=0.1M$ and $a=0.8M$, and several values of $jM^{2}$, excluding the region behind the event horizon. For sufficiently small $j$, we observe three disconnected surfaces: one compact region inherited from the black hole which encompasses the horizon, and two non-compact regions which extend to spatial infinity above and below the equatorial plane. The latter are already present in the swirling universe [4, 6]. The ergoregion in the upper hemisphere counter-rotates with respect to the black hole’s spin, i.e., $\Omega<0$ (coloured red), while the one in the lower hemisphere co-rotates with $\Omega>0$ (coloured blue). Note that the ergoregion inherited from the Kerr spacetime around the event horizon is always present for a finite value of the angular momentum parameter, but only becomes visible in the plots for large enough $a$ (see Fig. 2(b)).
@@ 3.1 Light rings
In the SBHSU, there is one degenerate LR on the equatorial plane at $r=3M$, when $j=0$, since in that case there is no distinction between co- and counter-rotation with respect to the swirling of the spacetime. However, as $j$ increases, the degeneracy is broken and the LRs branch off: one into the upper hemisphere and one into the lower as seen in Fig. 4(a). This pair of LRs is pushed off the equatorial plane and at the same time pulled towards the horizon. Note, in particular, that both LRs have the same radius. As $j\rightarrow\infty$, the LRs eventually reach the horizon $r=2M$, with one LR at the North pole $\theta=0$ and the other at the South pole $\theta=\pi$. In fact, for the SBHSU there is a symmetry of the effective potential such that if $(r,\theta)$ is a critical point of $H_{+}$ (shown in blue), then $(r,\pi-\theta)$ is also a critical point of $H_{-}$ (shown in orange) [16].
For the KBHSU where $a\neq 0$, this symmetry no longer holds, though for small $a$, we observe similar features as those appearing in the SBHSU case: the LRs are again pushed off the equatorial plane and pulled towards the horizon. They now, however, start from the two different branches of LR locations for vanishing $j$, since co- and counter-rotating rings no longer agree for Kerr black holes.
The spin-spin interaction of the KBHSU gives rise to an interesting new feature : the emergence of three coexisting LRs. We observe this effect for all the values of $a$ that we have studied, as demonstrated in Figs. 4(b), 4(c), 4(d), and 4(e). As seen in the figures, the third LR appears at the critical value $ajM=0.25$. The new LR is a critical point of $H_{-}$ which emerges from the axis $\theta=0$ at large $r$ and is then pulled inward while moving slowly away from the close vicinity of the axis. Shortly after the appearance of the third LR, however, the LR corresponding to the critical point of $H_{+}$ disappears.
@@
Figure 4: Locations of LRs at multiples of $jM^{2}=0.002$ with $M=1$. In the left-hand column we plot the coordinates of the LRs in the $(r,\theta)$-plane, in the middle column we give the radial coordinate against $j$, and in the right-hand column we show the polar coordinate against $j$. LRs corresponding to $H_{+}$ are coloured blue, and LRs corresponding to $H_{-}$ are in orange. The thick black dashed line represents the radius of the outer horizon. The pink vertical strip highlights the region where three LRs occur for the same value of $j$. The third LR appears around $ajM=0.25$ and always begins at the North pole $\theta=0$. The colours of the curves indicate whether the LR is a root of $H_{+}$ (light blue) or $H_{-}$ (orange).
In Fig. 6, we show the ergoregions for $a=0.5M$ and $a=0.9M$ and different values of $jM^{2}$ in the $(r,\theta)$-plane and include the locations of all of the LRs for the respective parameter sets as well as the ergoregions (highlighted in light blue). Here, we indicate whether the direction of rotation of the LR is prograde (blue) or retrograde (red) with respect to the angular momentum of the BH. To find the rotational direction of the photon rings, we consider the azimuthal rotation given by the quantity $\Phi$ which is defined as :
@@
$\Phi_{\pm}=H_{\pm}=\omega\pm\mathcal{F}\rho,$ (14)
and thus the rotation of the LR associated with the $H_{+}$ depends on the sign of $\Phi_{+}$ and the LR associated with the $H_{-}$ depends on the sign of $\Phi_{-}$. For the pure Kerr case, we always have two LRs on the equatorial plane: one retrograde orbit (red) and one prograde orbit (blue) which lies at a radius smaller than that of the retrograde orbit. This can be explained by the Lense-Thirring effect [19]. For sufficiently large BH angular momentum $a$, the prograde orbit lies inside the ergoregion (Fig. 6(b)). As $j$ increases and the ergoregions expand, both LRs are pulled off the equatorial plane and eventually lie within the ergoregions, thus co-rotating with the swirling background.

### v2

@@ Abstract
We discuss photon rings around as well as shadows of Kerr black holes immersed in a swirling spacetime (KBHSU). We find that the spin-spin interaction between the angular momentum of the black hole and the swirling of the background leads to new interesting effects as it breaks the symmetry between the upper and lower hemispheres. We find that a pair of light rings exists for all values of the parameter space. Using a topological argument, we prove that there should be, indeed, two light rings and that, additionally, these light rings are unstable. In comparison to the Schwarzschild black hole immersed in a swirling universe, the light rings typically all possess different radii. Interestingly, as the value of the swirling parameter is increased at fixed angular momentum of the black hole the two disconnected patches of the ergoregions eventually merge. The light ring at this merger possesses no angular velocity (as measured by an observer at infinity) and is called a light point. To our knowledge, this is the first time the existence of such a light point in a black hole space-time is reported. Finally, we also present the shadows of KBHSU for various parameter values and observe that, due to the presence of the swirling background, the shadows are twisted.
@@ 1 Introduction
Recently, a new exact, rotating solution has been constructed using the Ernst formalism [4]. This solution describes a rotating Kerr black hole in a swirling universe [5, 6]. The swirling universe, i.e. the solution without the black hole, is a vacuum solution equipped with an odd-$\mathbb{Z}_{2}$ symmetry: the upper ($\theta\in[0:\pi/2[$ ) and lower hemispheres ($\theta\in\ ]\pi/2:\pi]$) of the spacetime rotate in opposite directions. The swirling universe can, in fact, be obtained by using the Lie point symmetry through the application of the Ehlers transformation. Note that the geodesic motion can be decoupled and completely integrated in this spacetime [7]. An interesting question is whether this solution is of potential observational relevance, and which new features appear that allow distinction from the “pure” Kerr spacetime. An analysis of accretion discs in this spacetime has already been performed, revealing their bowl-like structure [8]. Recently, the interest in generating techniques has grown again, and different new solutions including the swirling background have been constructed. In particular, external electromagnetic fields were considered in a composition of Harrison and Ehlers’ transformation [9, 10], as well as charged black holes, i.e. Kerr-Newman black holes immersed in the swirling background [11, 9, 12]. Furthermore, new solutions can be obtained by exploring different symmetries. Recently, the inversion transformation has been used to produce the Kerr-Levi-Civita spacetime [13].
@@ 2 The spacetime and its properties
The metric tensor describing the KBHSU can be generated by applying an Ehlers transformation on a Kerr black hole seed solution within the Ernst formalism [4]. In Boyer-Lindquist coordinates, it reads  [^1] This is the metric as given in the auxiliary files provided with [4]. :
@@
$\omega_{0}$ $=$ $2aMr,\ \ \ \quad\omega_{1}=-4\cos\theta\left((r^{3}-a^{2}M)\Delta+a(r-M)\mathcal{W}\right)$
$\omega_{2}$ $=$ $2M\left[3ar^{5}-a^{5}(r+2M)+2a^{3}r^{2}(r+3M)-r^{3}(\cos^{2}\theta-6)\mathcal{W}+a^{2}\mathcal{W}\left(\cos^{2}\theta(3r-2M)+6(r-M)\right)\right]$ (3)
@@ 2.1 Conical singularity
KBHSU spacetime is non-asymptotically flat and has a non-removable conical singularity. We can identify the existence of an axial angular deficit/excess by considering the ratio between the perimeter and radius of a small circle around the rotational axis at both the north and the south poles [4, 9, 12]
@@
for non-vanishing angular momentum parameter $a$ and finite $j$. Thus the SBHSU does not suffer from conical singularities. We note the divergence for the critical value $ajM=\pm 0.25$. To understand whether this divergence corresponds to a physical singularity we compute the Kretschmann scalar $K=R_{\mu\nu\sigma\gamma}R^{\mu\nu\sigma\gamma}$ of the solution when $j=\pm 1/4aM$. The expressions are too lengthy and convoluted to be shown explicitly, but have the following form:
$\left.K\right|_{j=\pm 1/4aM}=\frac{\mathcal{G}\left(r,\theta\right)}{\left(1\mp\cos\theta\right)^{6}}\,.$ (6)
Therefore, the solution is singular along the upper semiaxis, $\theta=0$, for $4ajM=1$, and along the lower semiaxis, $\theta=\pi$, for 4ajM=14ajM=-1.
@@ 2.2 Ergoregions
A typical characteristic observed in the swirling family of solutions is the appearance of non-compact ergoregions extending from the equatorial plane to infinity [4, 8, 9, 12]. An ergoregion is defined by $g_{tt}>0$ with boundary at $g_{tt}=0$. For the KBHSU spacetime this reads $\omega^{2}/{\cal F}-\rho^{2}{\cal F}>0$ with boundary at $\omega^{2}=\rho^{2}{\cal F}^{2}$. The angular velocity $\Omega$ of the frame dragging is given by $\Omega=-g_{t\varphi}/g_{\varphi\varphi}=\omega$, which is not constant. We focus our analysis on the region outside the event horizon. The cross section of the ergoregions with the plane $y=0$ ($\varphi=0,\pi$) and their frame dragging directions are shown in Fig. 2 for two values of the angular momentum parameter, $a=0.1M$ and $a=0.8M$, and several values of $jM^{2}$, excluding the region behind the event horizon. For sufficiently small $j$, we observe three disconnected surfaces: one compact region inherited from the black hole which encompasses the horizon, and two non-compact regions which extend to spatial infinity above and below the equatorial plane. The latter are already present in the swirling universe [4, 7]. The ergoregion in the upper hemisphere counter-rotates with respect to the black hole’s spin, i.e., $\Omega<0$ (coloured red), while the one in the lower hemisphere co-rotates with $\Omega>0$ (coloured blue). Note that the ergoregion inherited from the Kerr spacetime around the event horizon is always present for a finite value of the angular momentum parameter $a$, but only becomes visible in the plots for large enough $a$ (see Fig. 2(b)).
@@ 3.1 Light rings
In the SBHSU (where $a=0$), for $j=0$ there is one degenerate LR on the equatorial plane at $r=3M$ since in that case there is no distinction between co- and counter-rotation with respect to the swirling of the spacetime. However, as $j$ increases, the degeneracy is broken and the LRs branch off: one into the upper hemisphere and one into the lower as seen in Fig. 4(a). This pair of LRs is pushed off the equatorial plane and at the same time pulled towards the horizon. Note, in particular, that both LRs have the same radius. As $j\rightarrow\infty$, the LRs eventually reach the horizon $r=2M$, with one LR at the North pole $\theta=0$ and the other at the South pole $\theta=\pi$. In fact, for the SBHSU there is a symmetry of the effective potential such that if $(r,\theta)$ is a critical point of $H_{+}$ (shown in blue), then $(r,\pi-\theta)$ is also a critical point of $H_{-}$ (shown in orange) [20].
For the KBHSU where $a\neq 0$, this symmetry no longer holds, though for small $a$, we observe similar features as those appearing in the SBHSU case: the LRs are again pushed off the equatorial plane and pulled towards the horizon. The branches now, however, start from the two different LR locations for vanishing $j$, since co- and counter-rotating rings no longer agree for Kerr black holes (see the left hand graphs in Figs. 4(b)-4(e)).
Due to the spin-spin interaction of the KBHSU, we observe that as $j$ increases to the critical value where $ajM=0.25$, the LR corresponding to the $H_{+}$ potential reaches the North pole $\theta=0$, and then disappears, passing over to $H_{-}$ and drawing away from the axis. We observe this effect for all the values of $a$ that we have studied, as demonstrated in Figs. 4(b)-4(e).
@@
Figure 4: Locations of the Light Rings for the KBHSU solution at multiples of $jM^{2}=0.002$ with $M=1$. LRs corresponding to $H_{+}$ are coloured blue and LRs corresponding to $H_{-}$ are in orange. In the left-hand column we plot the coordinates of the LRs in the $(r,\theta)$-plane, in the middle column we give the radial coordinate against $j$, and in the right-hand column we show the polar coordinate against $j$. The thick black dashed line represents the radius of the outer horizon.
In Fig. 6, we show the ergoregions for $a=0.5M$ and $a=0.9M$ and different values of $jM^{2}$ in the $(r,\theta)$-plane. We include the locations of the LRs for the respective parameter sets as well as the ergoregions (highlighted in light blue). Here, we indicate whether the direction of rotation of the LR is prograde (circles) or retrograde (squares) with respect to the angular momentum of the BH. To find the rotational direction of the photon rings, we consider the azimuthal rotation given by the quantity $\Phi$ which is defined as:
@@
$\Phi_{\pm}=\omega\pm\mathcal{F}\rho.$ (16)
Notice that this is identical to the potential $H_{\pm}$ (14), thus the rotation of the LR associated with the $H_{+}$ depends on the sign of $\Phi_{+}$ and the one with the $H_{-}$ depends on the sign of $\Phi_{-}$. For the pure Kerr case, we always have two LRs on the equatorial plane: one retrograde orbit and one prograde orbit which lies at a radius smaller than that of the retrograde orbit. This can be explained by the Lense-Thirring effect [26]. For sufficiently large BH angular momentum $a$, the prograde orbit lies inside the ergoregion (Fig. 6(b)). As $j$ increases and the ergoregions expand, both LRs are pulled off the equatorial plane and eventually lie within the ergoregions, thus co-rotating with the swirling background.